A capacitor consists of two similar square aluminium plates, each 10 cm × 10 cm mounted parallel and opposite to each other. What is their capacitance in μ F when distance between them is 1 cm and the dielectric is air


1
Expert's answer
2021-04-21T07:04:01-0400

The capacity of a flat capacitor is "C=\cfrac{\epsilon \cdot\epsilon_0 \cdot S}{d}"

Where

"\epsilon=1"

"\epsilon_0=8.85 \cdot 10^{-12} \cfrac{F}{m}"

"S=0.1 \cdot 0.1 m^2=1 \cdot 10^{-2} m^2"

"d=1 \cdot 10^{-2} m"

Then

"C=\cfrac{1 \cdot8.85 \cdot 10^{-12} \cdot 1 \cdot 10^{-2}}{1 \cdot 10^{-2}}=8.85 \cdot 10^{-12}F=0.00885\mu F"